Three numbers whose product is 64 are 2, 4,
and 8 because 2 x 4 x 8 = 64.
a. Find three numbers whose product is 2.4.
b. Find another set of three numbers whose
product is 2.4.
c. Find three numbers whose sum is 2.4.
d. Find another set of three numbers whose
sum is 2.4.

Answer by Edwin McCravy(20054)

Three numbers whose product is 64 are 2, 4,
and 8 because 2 x 4 x 8 = 64.
a. Find three numbers whose product is 2.4.
Just divide it by something and then divide that by something.

Start by dividing 2.4 by something, say 3.

2.4 ÷ 3 = 0.8

Now divide 0.8 by something, say 2

0.8 ÷ 2 = 0.4

So 0.4 × 3 × 2 = 2.4

You find another set of three numbers by dividing twice like that.

----------------------------------

c. Find three numbers whose sum is 2.4.
Just subtract something and then subtract something from what you get.

2.4 - 0.9 = 1.5

Subtract something from that:

1.5 - 0.7 = 0.8

So 0.8 + 0.7 + 0.9 = 2.4

-----------------------------

You find another set of three numbers by subtracting twice like that.

Edwin